Cutter Enterprises purchased equipment for $66,000 on January 1, 2018. The equipment is expected to have a five-year life and a residual value of $8,700.

Using the straight-line method, depreciation for 2019 and the equipment's book value at December 31, 2019, would be:

Multiple Choice

  • $11,460 and $34,380 respectively.

  • $11,460 and $43,080 respectively.

  • $13,200 and $52,800 respectively.

  • $26,400 and $39,600 respectively.
Answer #1

Annual depreciation=(Cost-Residual value)/Useful Life

=(66000-8700)/5

=$11460/year

Hence depreciation for 2019=$11460

Equipment book value at December 31,2019=Cost-Accumulated depreciation

=66000-(11460*2)=$43080

Hence the correct option is :

$11460 and $43080 respectively.